A solution contains 5.10×10-3 M sodium phosphate and 1.29×10-2 M ammonium carbonate.  
Solid calcium nitrate is added slowly to this mixture.

A. What is the formula of the substance that precipitates first?

formula =




B. What is the concentration of calcium ion when this precipitation first begins?  
[Ca2+] =  M
Answer #1

(A) It is possible to form two kinds of salt precipitation: calcium phosphate (Ca3(PO4)2) and calcium carbonate (CaCO3).


Ksp of Ca3(PO4)2 = 2.07×10 ^ -33


Ksp of calcium carbonate = 5×10 ^ -9


Therefore, the Ksp of Ca3(PO4)2 is smaller, so it will precipitate first.


Therefore, the formula = Ca3 (PO4) 2


(B) Ca3(PO4)2(s)<=> 3Ca2 +(aq)+ 2PO4 ^ 3-(aq)


Ksp = [3Ca2 +] ^ 3 [2PO4 ^ 3-] ^ 2


2.07×10 ^ -33 = [3Ca2 +] ^ 3 (2×5.10×10 ^ -3) ^ 2


[Ca2 +] = 9×10 ^ -11M


answered by: Gavin
Answer #2

A.


After mixing sodium phosphate and calcium nitrate, Ca3(PO4)2 precipitates are formed, with a Ksp value of 1 x 10-25.


Similarly, after mixing ammonium carbonate and calcium nitrate, CaCO3 precipitates are formed, with a Ksp value of 8.7 x 10-9.


Because the Ksp value of Ca3(PO4)2 is less than that of CaCO3, Ca3(PO4)2 will form precipitation first.


Therefore, the molecular formula of the first precipitated substance is Ca3(PO4)2

---------------------------


B.


The partial dissociation of Ca3(PO4)2 is:


Ca3(PO4)2(s)<-------------------> 3Ca2 +(aq)+ 2PO42-(aq)


The expression of the solubility product, that is, the Ksp of Ca3 (PO4) 2 (equal to the product of the molar concentration of the product at the equilibrium stage during the reaction process increases to increase its stoichiometric coefficient).


Ksp = [Ca2 +] 3. [PO43-] 2


[Ca2 +] 3 = Ksp / [PO43-] 2


[Ca2 +] 3 = 1 x 10-25 /(5.10×10-3) 2


[Ca2 +] 3 = 3.84 x 10-21


[Ca2 +] = (3.84 x 10-21) 1/3


[Ca2 +] = 1.57 x 10-7 M


Therefore, [Ca2 +] = 1.57 x 10-7 M


answered by: Gavin